# Cross-modal Retrieval Models for Stripped Binary Analysis

### v1

@@ 3.3 Model Training
Loss Function. We employ different loss functions for the two models due to their distinct architectures and training objectives, as shown in Figure 2. For the BinSeek-Embedding model, we utilize InfoNCE loss van den Oord et al. (2019) to optimize the cross-modal embedding space. Given a batch of $N$ positive pairs of NL descriptions and binary functions, we compute their embeddings using the BinSeek-Embedding model. The InfoNCE loss is defined as follows:
$\mathcal{L}_{e}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(q_{i},p_{i})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(q_{i},p_{j})/\tau)}$ (3)
where $q_{i}$ and $p_{i}$ are the embeddings of the NL description and pseudocode function in the $i$-th positive pair, respectively. The $\text{sim}(\cdot,\cdot)$ denotes the similarity function, which is cosine similarity in our practice, and $\tau$ is a temperature hyperparameter controlling the concentration level of the distribution.
@@
Training Setup. We train both models using the AdamW optimizer with a learning rate of 1e-4 and a cosine weight decay scheduler with minimum of 1e-5. For the BinSeek-Embedding model, we use an equivalent batch size of 5,120 and train for 2 epochs, with a maximum sequence length of 4,096 tokens. The BinSeek-Reranker model is trained with an equivalent batch size of 512 for 3 epochs, enlarging a maximum sequence length to 16,384 tokens for holding contextual information. Our experiments are conducted on H100 $\times$ 8 GPUs, each equipped with 80GB GPU memory, 2TB RAM, and Intel Xeon Platium 8468 with 48 cores. The whole training of BinSeek requires approximately 2k GPU hours in total.
@@ 4 Benchmarking Binary Code Retrieval
Since BinSeek incorporated two distinct models, we construct a comprehensive benchmark dataset to assess retrieval and reranking performance separately. We split a part of the raw dataset for test data construction, and ensure that the source code of these data does not overlap with training data to prevent data leakage. In total, we constructed 400 test data for embedding model evaluation and 200 for reranking model evaluation, and the details are as follows:
Evaluate Embedding Model. To evaluate the BinSeek-Embedding model’s retrieval capability, we create a search pool containing large-scale binary functions for each NL query, which is precisely how the model is applied in the real-world scenario. Specifically, for each positive pair $(q_{i},p_{i})$ in the synthesized corpus, where $q_{i}$ denotes the NL description and $p_{i}$ denotes the corresponding pseudocode function, we construct a search pool $\mathcal{P}_{i}$ of size $|\mathcal{P}_{i}|=K$ (empirically set to 10,000 in our experiments). The pool comprises the positive sample $p_{i}$ along with (K1)(K-1) carefully selected hard negative samples that are semantically unrelated to the positive one. This design ensures that no unlabeled positive sample leaks into the negative samples.
Based on our training data sampler (§3.1), we employ a filtering strategy to identify semantically unsimilar hard negatives. Given a candidate $p_{j}$ from the corpus, we assess its semantic relevance to the positive sample $p_{i}$ by comparing their associated NL descriptions, using a off-the-shelf embedding model (i.e., Qwen3-Embedding-8B). Notably, we further employ an LLM to generate rephrased semantic descriptions for the candidates, thereby obtaining a more comprehensive measure of correlation. Algorithm 1 of §A.3 comprehensively presents the processing for iteratively collecting test data with similarity-based filtering.
@@ 5.1 Evaluation Settings
Metrics. We employ two widely-adopted metrics to comprehensively measure our embedding and reranking models. Recall at $k$ (Rec@k) quantifies the proportion of queries for which the relevant item appears within the top-$k$ retrieved results, directly indicating the system’s success rate in retrieving the correct function among the top candidates. While, the Mean Reciprocal Rank at $k$ (MRR@k) further focuses on the rank position of the postive item, considering only the top-$k$ results for each query. MRR@k computes the average of the reciprocal ranks of the positive item, providing insight into how highly the correct function is ranked within the top candidates.
We report the metrics at $k=\{1,3,10\}$, which are chosen to reflect different practical scenarios in our retrieval task. Specifically, $k=1$ corresponds to the strictest setting, as each query is associated with a single positive instance; this setting evaluates the model’s precision in ranking the correct result at the very top. The choice of $k=3$ is motivated by practical tolerance for retrieval errors, where returning the top three results is considered acceptable. Finally, considering the whole retrieval-reranking system, we set the threshold $k=10$ to keep the initially retrieved candidates, who can be passed to the reranking model for further refinement.
Baselines. Notably, we are the first public work to explore NL-pseudocode retrieval model for stripped binary analysis, thus there is no existing work directly comparable to ours. For a deeper understanding of BinSeek’s performance, we conduct a comparison against existing general-purpose models. In selecting baselines, we consider factors including the SOTA performance in public leaderboards Muennighoff et al. (2022); Enevoldsen et al. (2025); Li et al. (2024), popularity among community, and accessibility for reproduction. Finally, we selected a total of 13 baseline models: 9 embedding models and 4 reranking models, detailed in Table 4 of §A.4.
@@ 5.2 Embedding & Reranking Evaluation
To answer RQ1, we conduct a comprehensive evaluation of our retrieval framework. We first assess the standalone retrieval performance of BinSeek-Embedding and the baselines. The test dataset has been outlined in §3.1, with each query containing one positive sample and 9,999 negative samples.
@@
First, we analyze the performance of BinSeek-Embedding on different metrics to understand its retrieval capability. Our model achieves 67% in Rec@1, which measures the ability to precisely identify the target function at the very first position among 10,000 candidates, demonstrating a strong discrimination capability in large-scale retrieval scenarios. The Rec@3 score indicates that our model can directly return useful results with an 80.5% probability in practical applications. More importantly, Rec@10 achieves 93.5%, suggesting that the vast majority of suboptimal retrieval results have the opportunity to be corrected to top-3 or even top-1 through subsequent reranking refinement. The MRR@10 of 75.18% further confirms that the correct results are typically ranked at high positions, demonstrating the semantic relevance of the retrieved results.
Second, we compare BinSeek-Embedding against baselines to highlight its advantages in both accuracy and efficiency. We first examine models with comparable parameter scale. Despite having a similar model size (0.3B), BinSeek-Embedding significantly outperforms embeddinggemma-300m, achieving 26.5% higher in Rec@1 and 28.53% higher in MRR@10. This substantial gap demonstrates that our model with domain-specific training is far more effective than general-purpose models, even when they have approximate potential. More remarkably, compared to the best baseline SFR-Embedding-Mistral that has $23\times$ more parameters, BinSeek-Embedding still achieves 16% higher Rec@10 and 9.02% higher MRR@10.
Table 2: The evaluation results of reranking models.
Model Sz.(B) $Rec@1$ $Rec@3$ MRR@3M\!R\!R@3
jina-reranker-v3 0.6 33.50 63.50 46.83
bge-reranker-v2-m3 0.6 34.50 57.00 44.33
Qwen3-Reranker-0.6B 0.6 34.50 62.50 46.42
Qwen3-Reranker-8B 8 62.50 80.50 70.83
BinSeek-Reranker 0.6 61.50 83.00 70.50
As shown in Table 2, our BinSeek-Reranker achieved 61.5% in Rec@1, meaning that the model successfully promoted the correct function to the top position in most cases. The Rec@3 score of 83.0% indicates that users can locate the target function within the top-3 reranked results with high confidence. The MRR@3 of 70.50% further confirms that the correct results are typically ranked at the top positions after reranking.
Compared to general-purpose rerankers with the same parameter scale of 0.6B, BinSeek-Reranker demonstrates substantial advantages. It outperforms Qwen3-Reranker-0.6B by 27% in Rec@1 and 24.08% in MRR@3. On the other hand, BinSeek-Reranker achieved competitive performance compared to Qwen3-Reranker-8B ($13\times$ larger), with 1% and 0.33% lower in Rec@1 and MRR@3 respectively, but 2.5% leading in Rec@3. These results indicate that general-purpose rerankers, despite their strong performance on standard NLP benchmarks, struggle to capture the semantic differences from stripped binary code.
@@ 5.3 Overall Evaluation
As shown in Table 3, BinSeek with our tailored models achieved the best performance, recording 76.75% in Rec@1 and 84.5% in Rec@3. Driven by embeddinggemma-300m and Qwen3-Reranker-0.6B that have the same model size as ours, the pipeline achieved 30.7% lower in Rec@3 and 27.17% lower in MRR@3, suggesting a significant effectiveness of our domain-specific training. Since the SFR-Embedding-Mistral and Qwen3-Reranker-8B have demonstrated the best performance among the evaluated embedding and reranking models, respectively, we thus employed them to investigate the best performance that can be expected from combinations of non-tailored models. As shown in the results, this combination achieved impressive scores, yet still lagged behind BinSeek by 6.75% and 4.63% in Rec@3 and MRR@3, respectively. More importantly, considering the 20.93 minutes overhead for each query, which is 10 times longer than BinSeek, it is challenging to be applied in practical scenarios.
When working as a whole system, BinSeek achieved a Rec@1 score of 76.75%, significantly surpassing the standalone BinSeek-Embedding’s score of 67%. This demonstrates that our 2-stage workflow design produces highly usable retrieval results in practice. Meanwhile, the system achieved improvements of 7.42% and 6.33% over the initial retrieval in MRR@3 and MRR@10, respectively, further confirming the effectiveness of context-augmented reranking.
@@ 6 Conclusion
In this paper, we introduced BinSeek, a two-stage cross-modal retrieval framework for the analysis of stripped binary code. It incorporated BinSeek-Embedding to learn the semantic relevance between binary code and natural language, and BinSeek-Reranker to further refine the results with context augmentation. For training the expert models, we also proposed an LLM-based data synthesis pipeline to generate large-scale high-quality data pairs automatically. Our evaluation demonstrates that BinSeek achieves SOTA performance with 84.5% in Rec@3 and 80.25% in MRR@3, surpassing the best-performing general-purpose models by 6.75% and 4.62%, respectively, also demonstrating higher practical value with $10\times$ speedup.
@@ A.1 Raw Dataset Details
To build diverse binaries, we utilized the makepkg archlinux.org (2025) pipeline equipped with wrapped compilers. makepkg is a script to automate the building of packages released by Arch Linux. We wrapped the gcc and clang compilers to customize the compilation process with custom build flags, such as optimization levels and debug information. Although the compilation process is automated by scripts, it is still prone to errors, such as dependency issues or configuration mismatches. Therefore, we further integrated AutoCompiler Hu et al. (2025), which is an LLM agent-based automated compilation pipeline, to resolve build failures and expand our collection. We stored the compiled debugging information into a distinct file, which is used to link a stripped binary function to its source code.
@@ A.2 Data Synthesis Details
As mentioned in §3.1, we employ an advanced large language model to generate semantic descriptions and assess their quality. Specifically, we use DeepSeek-V3 DeepSeek-AI et al. (2025), a SOTA open-source LLM with 671B parameters, to perform both generation and quality evaluation tasks. Here we present the details of these prompts:
@@ 4. You need to write comments into a JSON format string, for example:
A.3 Dataset Construction Algorithms.
@@
Algorithm 1 outlines the construction of the embedding evaluation dataset. It employs an iterative sampling strategy with a similarity-based filter (MaxSim) to collect hard negative samples that are semantically distinct from the query, ensuring a robust assessment of retrieval capability.
@@
A.4 Baseline Model Specifications.
@@
For embedding models, we primarily draw from two well-known leaderboards: MMTEB Enevoldsen et al. (2025) and CoIR Li et al. (2024) (records as of December 2025). From MMTEB, we selected top-ranked general-purpose text embedding models including Qwen3-Embedding Zhang et al. (2025), Jina-Embeddings Günther et al. (2025), multilingual-e5-large Wang et al. (2024), and embeddinggemma Vera et al. (2025). From CoIR, which specifically evaluates code-related retrieval tasks, we selected specialized models including SFR-Embedding Liu et al. (2024); Meng et al. (2024), BGE-M3 Chen et al. (2024). As for reranking models, we selected four top-performing models from the MMTEB leaderboard Enevoldsen et al. (2025): Qwen3-Reranker Zhang et al. (2025), BGE-Reranker-v2-m3 Chen et al. (2024), and Jina-Reranker-v3 Wang et al. (2025a). Table 4 lists the detailed specifications of all baseline models, including parameter size, embedding dimension, and maximum context length, facilitating a fair and transparent comparison.

### v2

@@ 3.3 Model Training
Loss Function. We employ different loss functions for the two models due to their distinct architectures and training objectives, as shown in Figure 2. For the BinSeek-Embedding model, we utilize InfoNCE loss 2019InfoNCE to optimize the cross-modal embedding space. Given a batch of $N$ positive pairs of NL descriptions and binary functions, we compute their embeddings using the BinSeek-Embedding model. The InfoNCE loss is defined as follows:
$\mathcal{L}_{e}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(q_{i},p_{i})/\tau)}{\sum_{p_{j}\in\{p_{j}^{-}\}\cup\mathcal{P}}\exp(\text{sim}(q_{i},p_{j})/\tau)}$ (4)
where $q_{i}$ and $p_{i}$ are the embeddings of the NL description and pseudocode function in the $i$-th positive pair, respectively, while the $p_{j}^{-}$ is the hard negative. The $\text{sim}(\cdot,\cdot)$ denotes the similarity function, which is cosine similarity in our practice, and $\tau$ is a temperature hyperparameter controlling the concentration level of the distribution.
@@
Training Setup. We train both models using the AdamW optimizer with a learning rate of 1e-4 and a cosine weight decay scheduler with minimum of 1e-5. For the BinSeek-Embedding, we use an equivalent batch size of 5,120 and train for 2 epochs, with a maximum sequence length of 4,096 tokens. The BinSeek-Reranker is trained with an equivalent batch size of 512 for 3 epochs, enlarging a maximum sequence length to 16,384 tokens for holding contextual information. Our experiments are conducted on H100 $\times$ 8 GPUs, each equipped with 80GB GPU memory, 2TB RAM, and Intel Xeon Platium 8468 with 48 cores. The whole training of BinSeek requires approximately 2k GPU hours in total.
@@ 4 Benchmarking Binary Code Retrieval
Since BinSeek incorporated two distinct models, we build a comprehensive benchmark dataset to assess retrieval and reranking performance separately. We split 10% of the raw dataset for test data construction, and ensure that the source code of these data does not overlap with training data to prevent data leakage. In total, we constructed 400 test data for embedding evaluation and 400 for reranking evaluation, and the details are as follows:
Evaluate Embedding Model. To evaluate the BinSeek-Embedding’s retrieval capability, we created a search pool containing large-scale binary functions for each NL query, mirroring its intended real-world application. Specifically, for each positive pair $(q_{i},p_{i})$ in the synthesized corpus, where $q_{i}$ denotes the NL description and $p_{i}$ denotes the corresponding pseudocode function, we constructed a search pool $\mathcal{P}_{i}$ of size $|\mathcal{P}_{i}|=K$ (empirically set to 10,000 in our experiments). The pool comprises the positive sample $p_{i}$ along with (K1)(K-1) carefully selected negative samples that are semantically unrelated to the positive one.
Given a candidate $p_{j}$ from the corpus, we assess its semantic relevance to the positive sample $p_{i}$ by comparing their associated NL descriptions. Following our training data sampler (§3.1), we employed the same threshold of similarity score for filtering. This design ensures that no unlabeled positive sample leaks into the negative samples. Notably, we further employ an LLM to generate rephrased semantic descriptions for the candidates, thereby obtaining a more comprehensive measure of correlation. Algorithm 1 of §A.4 comprehensively presents the processing for iteratively collecting test data with similarity-based filtering.
@@
In summary, we sampled a moderately sized test dataset from a large-scale test corpus, with the primary goal of ensuring practical usability and ease of adoption. To verify its representativeness, we reconstructed the test set via repeated resampling from the raw corpus, observing a standard deviation of 1.52% in our evaluation metric across trials. Furthermore, repeated subsampling of the fixed final test queries yields a lower standard deviation of 0.91%, indicating stable evaluation results. Together, these results suggest that our test dataset achieves a favorable balance between usability, representativeness, and statistical stability.
@@ 5.1 Evaluation Settings
Metrics. We employ two widely-adopted metrics to comprehensively measure the embedding and reranking models. Recall at $k$ (Rec@k) quantifies the proportion of queries for which the relevant item appears within the top-$k$ retrieved results, directly indicating the system’s success rate in retrieving the correct function among the top candidates. While, the Mean Reciprocal Rank at $k$ (MRR@k) further focuses on the rank position of the postive item, considering only the top-$k$ results for each query. MRR@k computes the average of the reciprocal ranks of the positive item, providing insight into how highly the correct function is ranked within the top candidates.
We report the metrics at $k=\{1,3,10\}$, which are chosen to reflect different practical scenarios in our retrieval task. Specifically, $k=1$ corresponds to the strictest setting, as each query is associated with a single positive instance; this setting evaluates the model’s precision in ranking the correct result at the very top. The choice of $k=3$ is motivated by practical tolerance for retrieval errors, where returning the top three results is considered acceptable. Finally, considering the whole retrieval-reranking system, we set $k=10$ to keep the initially retrieved candidates, who can be passed to the reranking model for further refinement.
Baselines. To the best of our knowledge, the only related work is BinQuery binquery, which, however, does not publicly release their model weights and dataset, making direct comparison infeasible. Moreover, BinQuery focuses on disassembly code, which differs from our NL-pseudocode retrieval task. Therefore, there are no publicly available domain-specific methods that can be directly compared with ours. As a result, we conduct a comparison against existing general-purpose models. In selecting baselines, we consider factors including the SOTA performance in public leaderboards 2022mteb; 2025mmteb; CoIR, popularity among community, and accessibility for reproduction. Finally, we selected a total of 13 baseline models: 9 embedding models and 4 reranking models, detailed in Table 4 of §A.5.
@@ 5.2 Embedding & Reranking Evaluation
To answer RQ1, we conducted a comprehensive evaluation of our retrieval framework. We first assess the standalone retrieval performance of BinSeek-Embedding and the baselines. The test dataset has been outlined in §3.1, with each query containing one positive and 9,999 negatives.
@@
First, we analyze the performance of BinSeek-Embedding on different metrics to understand its retrieval capability. The Rec@1 score of 67% demonstrates essentially strong performance in large-scale retrieval. The Rec@3 score indicates that our model can directly return useful results with an 80.5% probability in practical acceptance. More importantly, Rec@10 achieves 93.5%, suggesting that the vast majority of suboptimal results have the opportunity to be corrected to top-3 or even top-1 through subsequent reranking. The MRR@10 of 75.18% further confirms that the correct results are typically ranked at high positions, demonstrating the semantic relevance of retrieved results.
@@
Second, we compare BinSeek-Embedding with baselines to highlight its advantages in both accuracy and efficiency. We first examine models with similar parameter scale. Despite having a close size (0.3B), BinSeek-Embedding significantly outperforms embeddinggemma-300m, achieving 26.5% higher in Rec@1 and 28.53% higher in MRR@10. More remarkably, compared to the best baseline SFR-Embedding-Mistral that has $23\times$ more parameters, BinSeek-Embedding still achieves 16% higher Rec@10 and 9% higher MRR@10.
Table 3: The evaluation results of reranking models.
Model Sz.(B) $Rec@1$ $Rec@3$ MRR@3M\!R\!R@3
jina-reranker-v3 0.6 34.00 63.75 47.33
bge-reranker-v2-m3 0.6 34.00 57.75 46.03
Qwen3-Reranker-0.6B 0.6 35.00 62.00 46.42
Qwen3-Reranker-8B 8 62.75 81.00 70.94
BinSeek-Reranker 0.6 61.75 83.00 70.23
As shown in Table 3, our BinSeek-Reranker achieved 61.75% in Rec@1, meaning that the model successfully promoted the correct function to the top position in most cases. The Rec@3 score of 83,0% indicates that users can locate the target function within the top-3 reranked results with high confidence. While the MRR@3 of 70.23% further confirms that the correct results are typically ranked at the top positions after reranking.
Compared to general-purpose rerankers with the same parameter scale of 0.6B, BinSeek-Reranker demonstrates substantial advantages. It outperforms Qwen3-Reranker-0.6B by 27.75% of Rec@1 and 22.9% of MRR@3. Moreover, BinSeek-Reranker achieved competitive performance compared to Qwen3-Reranker-8B ($13\times$ larger), with 1% and 0.71% lower in Rec@1 and MRR@3 respectively, but 2% leading in Rec@3. These results demonstrate that our domain training constructed effective and efficient expertise reranker.
Additionally, our context-aware design employed a heuristic informative measurement to select context functions as augmented information for improving reranking performance. Experiments demonstrate the effectiveness of this design, boosting our model and the baselines by 8% and 4% Rec@1, respectively. Details can be found in §A.6.
@@ 5.3 Overall Evaluation
As shown in Table 2, BinSeek with our tailored models achieved the best performance, recording 76.75% in Rec@1 and 84.5% in Rec@3. Driven by embeddinggemma-300m and Qwen3-Reranker-0.6B that have the same model size as ours, the pipeline achieved 30.7% lower in Rec@3 and 27.17% lower in MRR@3, suggesting a significant effectiveness of our domain-specific training. Since the SFR-Embedding-Mistral and Qwen3-Reranker-8B have demonstrated the best performance among the evaluated embedding and reranking baselines, respectively, we thus employed them to investigate the best performance that can be expected from non-tailored models. As shown in the results, this combination achieved impressive scores, yet still lagged behind BinSeek by 6.75% and 4.63% in Rec@3 and MRR@3, respectively. More importantly, considering the 20.93 minutes overhead for each query, which is 10 times longer than BinSeek, it is challenging to be applied in practical scenarios. When working as a whole system, BinSeek’s Rec@1 score is 9.75% higher than 67% of the BinSeek-Embedding standalone. This demonstrated that our 2-stage workflow design produces highly usable retrieval in practice.
@@ 6 Conclusion
In this paper, we introduced BinSeek, a two-stage cross-modal retrieval framework for the analysis of stripped binary code. It incorporated BinSeek-Embedding to learn the semantic relevance between binary pseudocode and NL, and BinSeek-Reranker to further refine the results with context augmentation. For training our models, we also proposed an LLM-based data synthesis pipeline to generate large-scale high-quality data pairs automatically. Our evaluation demonstrated that BinSeek achieved SOTA performance with 84.5% in Rec@3 and 80.25% in MRR@3, surpassing the best-performing general-purpose models by 6.75% and 4.62%, respectively, also demonstrating higher practical value with $10\times$ faster speed.
@@ 7 Limitations
8 Ethical Considerations
While BinSeek is designed to support defensive security workflows (e.g., vulnerability investigation, program auditing, and malware analysis by defenders), it may also lower the barrier to reverse engineering and be misused by malicious actors. In particular, by enabling natural language search over stripped binary codebases, our retrieval models could help attackers more quickly understand a target binary, locate security relevant functionality, and accelerate the analysis and development of malicious behaviors. Therefore, this technology can benefit legitimate security research and defense, but it also carries a risk of misuse.
@@ A.1 Raw Dataset Details
To build diverse binaries, we utilized the makepkg makepkg pipeline equipped with wrapped compilers. makepkg is a script to automate the building of packages released by Arch Linux. We wrapped the gcc and clang compilers to customize the compilation process with custom build flags, such as optimization levels and debug information. Although the compilation process is automated by scripts, it is still prone to errors, such as dependency issues or configuration mismatches. Therefore, we further integrated AutoCompiler ACL2025compileagent, which is an LLM agent-based automated compilation pipeline, to resolve build failures and expand our collection. We stored the compiled debugging information into a distinct file, which is used to link a stripped binary function to its source code.
@@ A.2 Data Synthesis Details
As mentioned in §3.1, we employ an advanced large language model to generate semantic descriptions and assess their quality. Specifically, we use DeepSeek-V3 2025deepseekv3, a SOTA open-source LLM with 671B parameters, to perform both generation and quality evaluation tasks. Here we present the details of these prompts:
@@ 4. You need to write comments into a JSON format string, for example:
A.3 Human Evaluation on Synthetic Labels
To validate the quality of the LLM-generated natural language descriptions used for training, we conducted a human evaluation on a randomly sampled subset of the synthesized dataset. We randomly selected 100 function–description pairs. For each sample, human evaluators were provided with the original source-level function along with the same meta information that was available to the LLM during label generation, while the LLM discriminator’s assessment was not shown.
Three human evaluators independently participated in the evaluation: one reverse engineering expert with eight years of industry experience, and two graduate students with four and three years of experience in binary analysis, respectively. To mitigate potential fatigue and mechanical judgments, each evaluator assessed 20 samples per day, spending about 30 minutes per session.
Evaluators were asked to judge “whether the generated description accurately reflects the core functionality of the corresponding function, using three categories: Correct, Partially correct, or Incorrect”. Across all evaluators, an average of 90.3% of the generated descriptions were rated as Correct, and the percentage increases to 97.6% when including Partially correct cases. This indicates that the LLM-generated labels are largely accurate and provide reliable supervision for model training.
A.4 Dataset Construction Algorithms.
@@ A.3 Human Evaluation on Synthetic Labels
Algorithm 1 outlines the construction of the embedding evaluation dataset. It employs an iterative sampling strategy with a similarity-based filter (MaxSim) to collect real negative samples that are semantically distinct from the query, ensuring a robust assessment of retrieval capability.
@@
A.5 Baseline Model Specifications.
@@
For embedding models, we primarily draw from two well-known leaderboards: MMTEB 2025mmteb and CoIR CoIR (records as of December 2025). From MMTEB, we selected top-ranked general-purpose text embedding models including Qwen3-Embedding 2025qwen3-emb, Jina-Embeddings 2025jinaembeddingsv4, multilingual-e5-large wang2024multilingual, and embeddinggemma vera2025embeddinggemma. From CoIR, which specifically evaluates code-related retrieval tasks, we selected specialized models including SFR-Embedding liu2024codexembed; 2024SFR, BGE-M3 2024bge-m3. As for reranking models, we selected four top-performing models from the MMTEB leaderboard 2025mmteb: Qwen3-Reranker 2025qwen3-emb, BGE-Reranker-v2-m3 2024bge-m3, and Jina-Reranker-v3 wang2025jinarerankerv3. Table 4 lists the detailed specifications of all baseline models, including parameter size, embedding dimension, and maximum context length, facilitating a fair and transparent comparison.
A.6 Ablation Studies on Binary Code Context
Figure 6: Comparison of the selection method of binary code context across different metrics and rerankers. “ctx” means the context functions are selected by our heuristic informative score $\mathcal{S}$, and “rndctx” means randomly selecting context functions. While “w/o ctx” means use no context in reranking.
To further investigate the effectiveness of context augmentation and our heuristic informative measurement, we conducted an ablation study on the context selection method. We compared the performance of the following three methods: (1) using our heuristic informative measurement to select the context functions with top-$5$ scored, (2) randomly selecting $5$ context functions, and (3) using no context in reranking. We involved our BinSeek-Reranker and a representative baseline model Qwen3-Reranker-8B in this experiment.
As shown in Figure 6, both BinSeek-Reranker and Qwen3-Reranker-8B achieved the best performance when using our heuristic informative measurement to select the context functions. In particular, the selected context boosted our model and the baseline model with 8% and 4% Rec@1, respectively. Meanwhile, randomly selecting context functions also achieved better performance than using no context. This demonstrates the effectiveness of our context-aware design. On the other hand, randomly selected context achieved suboptimal performance compared to employing our selection method, suggesting the contribution of our informative measurement.